OpenStudy (anonymous):

x^2-2x-y-8=0 I need to find the standard form of the equation of this parabola.

jhonyy9 (jhonyy9):

y=x2-2x-8

jhonyy9 (jhonyy9):

x2 is x on exponent 2

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

That is not one of the 4 possible answers given,Must be (x plus or minus 1)squared then plus or minus 9... i.e. (x-1)^2-9

OpenStudy (anonymous):

that is y= ( x-1)^2-9....?

jhonyy9 (jhonyy9):

y=(x-4)(x+2)=x2 -2x -8

OpenStudy (anonymous):

I see what you are saying, but it must be, (x-1)^2-9 (x-1)^2+9 (x+1)^2-9 (x+1)^2+9 these are the four given answers in the book.

jhonyy9 (jhonyy9):

ok